Velocity of air is 320 m/s a pipe is  closed at one end  has length of 1m neglecting end correction then the air column can resonate with a frequency of

Grade:Select Grade

2 Answers

Jay Sawant
6 years ago
Here the pipe is opened at one end and closed at other end.Therefore, at open end the displacement wave will always have antinode while at the closed end the displacement wave will always have node.During its fundamental freq , wavelength=4l(l=length of pipe)f = v/4lDuring 1st Overtone Wavelength=4/3lf = 3(v/4l)..Now doing the same for 2nd,3rd and so on overtonesFrequency=n(v/4l) n belongs to odd integers(n>0).These are called as odd harmonics.
Arun
6 years ago
For a closed pipe, frequrncies of harmonics are
\LARGE \nu1 = v/4L = 320/4 = 80 Hz
\LARGE \nu2 = 3v/4L = 3 *80 = 340 Hz
\LARGE \nu3 = 5v/4L = 5*80 = 400 Hz
Similarily we can also get some more frequencies